# LLM Unlearning with LLM Beliefs

### v1

@@ 4 New Method: Bootstrapping-based Unlearning
Building on our observations in Section 3, in this section, we motivate a belief-aware objective against the squeezing effect and instantiate a bootstrapping framework at token and sequence levels in Section 4.1, and theoretically explain how it mitigates squeezing effect in Section 4.2.
@@ 4.1 Algorithm: Bootstrapping at Token and Sequence Levels
Motivation: From the Squeezing Effect to Bootstrapping. Analyses in Section 3 show that suppressing the exact target does not remove underlying knowledge; instead, probability mass is squeezed into semantically proximate regions already favored by the model. Given a forget prompt ${\mathbf{x}}_{\rm u}$ and prefix ${\mathbf{y}}_{\rm u}^{<i}$, the conditional distribution $\pi_{{\bm{\theta}}}(\cdot\mid{\mathbf{x}}_{\rm u},{\mathbf{y}}_{\rm u}^{<i})$ captures the model’s local belief at position $i$. The high-likelihood neighborhood can be approximated by the top-$k$ set $\mathcal{H}^{(i)}_{k}=\mathrm{Top}\text{-}k(\pi_{{\bm{\theta}}}(\cdot\mid{\mathbf{x}}_{\rm u},{\mathbf{y}}_{\rm u}^{<i}))$. At the sequence level, high-confidence generations $\hat{{\mathbf{y}}}_{\rm u}\sim\pi_{{\bm{\theta}}}(\cdot\mid{\mathbf{x}}_{\rm u})$ with large average log-likelihood represent the model’s global beliefs. Empirically, while GA and NPO decrease $\pi_{{\bm{\theta}}}(y^{i}_{\rm u}\mid{\mathbf{x}}_{\rm u},{\mathbf{y}}_{\rm u}^{<i})$ for the labeled token, they simultaneously increase mass on $\mathcal{H}^{(i)}_{k}$, producing high-confidence rephrasings that preserve sensitive content. Thus, spurious unlearning arises not from metric artifacts but from normalization-driven alignment with internal beliefs.
@@
Bootstrapping-Sequence (BS-S). While BS-T addresses local beliefs at the token level, it cannot fully prevent harmful continuations from re-emerging in longer outputs. BS-S extends bootstrapping to the sequence level, targeting the model’s global beliefs. Concretely, for each forget prompt ${\mathbf{x}}_{\rm u}$, we sample $N$ high-confidence generations $\hat{{\mathbf{y}}}_{\rm u}^{(j)}\sim\pi_{{\bm{\theta}}}(\cdot|{\mathbf{x}}_{\rm u})$ using temperature-controlled decoding, and construct an augmented unlearning set $\hat{\mathcal{D}}_{\rm u}=\{({\mathbf{x}}_{\rm u},\hat{{\mathbf{y}}}_{\rm u}^{(j)})\}_{j=1}^{N}$. By including these high-likelihood continuations in the forget set, BS-S exposes deeper memorization and ensures that entire harmful trajectories are suppressed. The final objective is
@@
where $\lambda_{\rm BSS}$ balances forgetting of the original targets and their bootstrapped augmentations, and $\mathcal{L}$ can be instantiated by any unlearning loss such as $\mathcal{L}_{\rm GA}$ or $\mathcal{L}_{\rm BST}$.
BS-T and BS-S are compatible with existing unlearning objectives such as NPO and WGA, and can also integrate regularization like GradDiff. As shown in Section 5, both bring clear gains: BS-T offers higher efficiency, while BS-S achieves more thorough forgetting. Pseudocodes are provided in Appendix C.
4.2 Theoretical Analysis: How BS-T Mitigates Squeezing Effect?
We provide a theoretical analysis of why BS-T mitigates the squeezing effect. Our analysis builds on the learning dynamics framework of LLM finetuning (Ren & Sutherland, 2025), which formalizes how a single SGD step affects the log-likelihood of a candidate response by decomposing the update into three parts: a Jacobian capturing softmax normalization, a kernel similarity transporting influence across examples, and a residual term reflecting the direct impact of the loss. Lemma 4.1 restates this decomposition and highlights the residual as the key factor driving probability shifts, while Theorem 4.2 compares the residuals of GA and BS-T to reveal how BS-T distributes forgetting pressure over both targets and their belief neighborhoods. Detailed proofs are deferred to Appendix D.
Lemma 4.1 (AKG Decomposition (Ren & Sutherland, 2025)).
@@ 4.2 Theoretical Analysis: How BS-T Mitigates Squeezing Effect?
where $\mathcal{A}_{t}(\bm{\mathsf{\chi}}_{\rm o})=\bm{I}-\mathbf{1}\pi^{\top}_{{\bm{\theta}}^{t}}(\cdot|\bm{\mathsf{\chi}}_{\rm o})$ is the softmax Jacobian, $\mathcal{K}_{t}(\bm{\mathsf{\chi}}_{\rm o},\bm{\mathsf{\chi}}_{\rm u})=\nabla_{{\bm{\theta}}}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o})\nabla_{{\bm{\theta}}}^{\top}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm u})$ is the eNTK, and $\mathcal{G}_{t}(\bm{\mathsf{\chi}}_{\rm u})=\nabla_{{\mathbf{z}}}\mathcal{L}(\bm{\mathsf{\chi}}_{\rm u})$ captures the residual term induced solely by the unlearning loss. Here ${\mathbf{z}}=h_{{\bm{\theta}}}(\bm{\mathsf{\chi}})$ denotes the token–logit matrix and all quantities are evaluated at ${\bm{\theta}}^{t}$.
Lemma 4.1 indicates that the update is mainly governed by $\mathcal{G}$: it specifies which tokens are pushed down or up before being modulated by $\mathcal{A}$ and transported via $\mathcal{K}$. Therefore, distinguishing the different forgetting behaviors of GA and BS-T reduces to analyzing the formulation of their residuals.
Theorem 4.2 (Residual Terms for GA vs. BS-T).
Under Lemma 4.1, denote ${\mathbf{q}}^{i}=\mathrm{sg}\big[\pi_{{\bm{\theta}}^{t}}(\cdot|\bm{\mathsf{\chi}}_{\rm u})\big|_{\mathcal{H}^{(i)}_{k}}\big$], the residual terms $\mathcal{G}$ for GA and BS-T at position $i$ are: (1) For GA, $\mathcal{G}^{i}_{\rm GA}=\pi_{{\bm{\theta}}^{t}}(\cdot|\bm{\mathsf{\chi}}_{\rm u})-{\mathbf{e}}_{y_{\rm u}^{i}}$; (2) For BS-T, $\mathcal{G}^{i}_{\rm BST}=\pi_{{\bm{\theta}}^{t}}(\cdot|\bm{\mathsf{\chi}}_{\rm u})-\big((1-\lambda){\mathbf{e}}_{y_{\rm u}^{i}}+\lambda\,{\mathbf{q}}^{i}\big)$. Hence for any component $v\neq y_{\rm u}^{i}$, we have
@@
Remark. Figure 3 provides intuitive illustration for Theorem 4.2. In GA, the gray curve $\pi_{\bm{\theta}_{\mathrm{o}}}$ shows the distribution before unlearning and the green curve $\pi_{{\bm{\theta}_{\mathrm{u}}}}$ after unlearning: the residual $\mathcal{G}_{\rm GA}$ pushes down the target $y_{\rm u}$ but reallocates mass to nearby high-likelihood regions, leading to semantically similar rephrasings. In BS-T, the shaded area marks the top-$k$ belief ${\mathbf{q}}^{i}$, and the residual $\mathcal{G}_{\rm BST}$ distributes repulsion across both the target and its close alternatives. The resulting blue curve suppresses the whole neighborhood rather than creating a new peak, reducing rephrasings and enabling more generalizable unlearning.
Table 1: Performance of different unlearning methods with retain regularization on TOFU with LLaMA-3-1B/3B/8B under 1%/5%/10% forget settings.
LLaMA-3.2-1B LLaMA-3.2-3B LLaMA-3.1-8B
@@
  • Notes: Agg. is the harmonic mean of Mem. and Util.. Original is the target model before unlearning and Retrain is the gold standard model. ↑/↓ indicate larger/smaller values are preferable. The best and runner-up results are bolded and underlined.
@@ 5.1 Experimental Setup
Benchmarks and Baselines. We assess unlearning performance across three representative benchmarks: TOFU (Maini et al., 2024), MUSE (Shi et al., 2025), and WMDP (Li et al., 2024b). Our approach is compared with representative baselines from OpenUnlearning (Dorna et al., 2025) incorporating retain regularization, including GradDiff (Maini et al., 2024), NPO (Zhang et al., 2024), RMU (Li et al., 2024b), SimNPO (Fan et al., 2025), and WGA (Wang et al., 2025b).
Model Architectures. We adopt a variety of LLM families for unlearning, including LLaMA-2 (Touvron et al., 2023), LLaMA-3 (Grattafiori et al., 2024), and Zephyr (Tunstall et al., 2024). Specifically, for TOFU, we employ LLaMA-3.2-1B/3B-Instruct and LLaMA-3.1-8B-Instruct. For MUSE and WMDP, we use LLaMA-2-7B-chat and Zephyr-7B-beta, respectively.
@@
Training Configuration. Our experiments follow the setups in OpenUnlearning as well as Wang et al. (2025d); Yang et al. (2025). The AdamW optimizer (Loshchilov & Hutter, 2019) is employed across all benchmarks. On TOFU, we train for 10 epochs with batch size 32, learning rate 1e-5, linear scheduler with 1 warm-up epoch, and weight decay 0.01. On MUSE, we use batch size 32, learning rate 1e-5, a constant scheduler, and select results from 10 checkpoints across 10 epochs. On WMDP, we train for 125 steps with batch size 16, learning rate 4e-6, and 25 warm-up steps.
For further details and introductions of the experimental setup, please refer to Appendix E.
@@ 5.2 Experimental Results
Results on TOFU. Table 1 summarizes results on TOFU under 1%, 5%, and 10% forget settings. Across all model scales (LLaMA-3-1B/3B/8B), our bootstrapping methods achieve superior performance. In particular, BS-S consistently delivers the best aggregate and memorization scores (e.g., Agg. 0.58/0.60/0.60 at 5% and 0.57/0.50/0.49 at 1%), clearly surpassing NPO and RMU. BS-T also ranks second in most cases, balancing forgetting and retention—for instance, Agg. 0.55 at 5%–3B and 0.54 at 1%–1B—while retaining competitive utility with higher efficiency. These findings confirm that unlearning both targets and model beliefs enables BS-S to achieve the most thorough forgetting, with BS-T as a strong runner-up, validating the effectiveness of our framework on TOFU.
Table 2: Performance with retain regularization on MUSE-News with LLaMA-2-7B-chat.
Forget Retain
Method VerbMem KnowMem UtilPres
Original $0.5789$ $0.6443$ $0.5552$
Retrain $0.2016$ $0.3170$ $0.5602$
GradDiff $0.3249$ $0.3481$ $0.4603$
NPO $0.2914$ $0.3290$ $0.4651$
RMU $0.3861$ $0.5088$ $0.4962$
SimNPO $0.4608$ $0.6043$ $0.5010$
WGA $0.3713$ $0.5732$ $0.4782$
BS-T (Ours) $0.2837$ $0.3278$ $0.4602$
BS-S (Ours) $0.2713$ $0.3250$ $0.4774$
  • Notes: We bolden the best results and underline the runner-ups. ↑/↓ indicates larger/smaller values are preferable.
Results on MUSE. Table 2 reports results on MUSE-News with LLaMA-2-7B-chat. Our methods achieve the lowest forget scores, with BS-S reducing VerbMem/KnowMem to 0.2713/0.3250 compared to 0.2914/0.3290 for NPO and 0.3861/0.5088 for RMU. At the same time, BS-S attains higher UtilPres (0.4774) than most baselines, showing a better balance between forgetting and retention. BS-T also improves forgetting effectiveness (0.2837/0.3278) while maintaining competitive utility (0.4602). See Appendix F.3 for additional results on MUSE-Books.
Table 3: Performance with retain regularization on WMDP with Zephyr-7B-beta.
@@
  • Notes: ↑/↓ indicates larger/smaller values are preferable. The best and runner-up results are bolded and underlined, respectively.
Results on WMDP. Table 3 presents results on WMDP with Zephyr-7B-beta. Recall that the forget score corresponds to QA accuracy, where values closer to 0.25 indicate more randomized responses and thus stronger unlearning. Both BS-T and BS-S achieve lower scores on Bio (0.26) and Cyber (0.28/0.27) compared with NPO (0.27/0.30) and RMU (0.29/0.27), while also attaining higher MMLU retention (0.52 and 0.54 vs. 0.44–0.48 for most baselines) except for RMU (0.55). Overall, BS-S delivers the best trade-off, reaching near-random forgetting accuracy while preserving more utility than most competing methods.
Figure 4: LaaJ evaluation on TOFU 10%.
LaaJ Evaluation. We adopt Gemini 2.5 Flash (Comanici et al., 2025) to assess Naturalness and Similarity in Section 3.1 with LLaMA-3-8B on TOFU 10%. As shown in Figure 4, our methods strike a robust trade-off between the two dimensions. Notably, BS-T and BS-S yield higher Naturalness and Similarity scores than the baselines (where higher values are preferred). This result indicates that our framework mitigates spurious unlearning, producing responses that remain fluent while being sufficiently divergent.
Additional Results in Appendix F. In addition to the content already been mentioned above, Appendix F.4 provides qualitative comparisons of unlearned responses across different unlearning methods; Appendix F.5 presents ablation studies covering hyperparameter analysis and the influence of different unlearning losses in BS-S; and Appendix F.6 reports training time comparisons.

### v2

@@ 4 New Method: Bootstrapping-based Unlearning
Building on our observations in §3, in this section, we motivate a belief-aware objective against the squeezing effect in §4.1 and instantiate a bootstrapping-based unlearning framework in §4.2.
4.1 Motivation: From the Squeezing Effect to Bootstrapping
Analyses in §3 show that suppressing the exact target does not remove underlying knowledge; instead, probability mass is squeezed into semantically proximate regions already favored by the model. Given a forget prompt ${\mathbf{x}}_{\rm u}$ and prefix ${\mathbf{y}}_{\rm u}^{<i}$, the conditional distribution $\pi_{{\bm{\theta}}}(\cdot\mid{\mathbf{x}}_{\rm u},{\mathbf{y}}_{\rm u}^{<i})$ captures the model’s local belief at position $i$. The high-likelihood neighborhood can be approximated by the top-$k$ set $\mathcal{H}^{(i)}_{k}=\mathrm{Top}\text{-}k(\pi_{{\bm{\theta}}}(\cdot\mid{\mathbf{x}}_{\rm u},{\mathbf{y}}_{\rm u}^{<i}))$. At the sequence level, high-confidence generations $\hat{{\mathbf{y}}}_{\rm u}\sim\pi_{{\bm{\theta}}}(\cdot\mid{\mathbf{x}}_{\rm u})$ with large average log-likelihood represent the model’s global beliefs. Empirically, while GA and NPO decrease $\pi_{{\bm{\theta}}}(y^{i}_{\rm u}\mid{\mathbf{x}}_{\rm u},{\mathbf{y}}_{\rm u}^{<i})$ for the labeled token, they simultaneously increase mass on $\mathcal{H}^{(i)}_{k}$, producing high-confidence rephrasings that preserve sensitive content. Thus, spurious unlearning arises not from metric artifacts but from normalization-driven alignment with internal beliefs.
@@ 4.2 Algorithm: Bootstrapping at Token and Sequence Levels
Bootstrapping-Sequence (BS-S). While BS-T addresses local beliefs at the token level, it cannot fully prevent harmful continuations from re-emerging in longer outputs. BS-S extends bootstrapping to the sequence level, targeting the model’s global beliefs. Concretely, for each forget prompt ${\mathbf{x}}_{\rm u}$, we sample $N$ high-confidence generations $\hat{{\mathbf{y}}}_{\rm u}^{(j)}\sim\pi_{{\bm{\theta}}}(\cdot|{\mathbf{x}}_{\rm u})$ using temperature-controlled decoding, and construct an auxiliary unlearning set $\hat{\mathcal{D}}_{\rm u}=\{({\mathbf{x}}_{\rm u},\hat{{\mathbf{y}}}_{\rm u}^{(j)})\}_{j=1}^{N}$. By including these high-likelihood continuations in the forget set, BS-S exposes deeper memorization and ensures that entire harmful trajectories are suppressed. The final objective is
@@
where $\lambda_{\rm BSS}$ balances forgetting of the original targets and their bootstrapped augmentations, and $\mathcal{L}$ can be instantiated by any unlearning loss such as $\mathcal{L}_{\rm GA}$ or $\mathcal{L}_{\rm BST}$. In practice, BS-S may operate in an off-policy form by sampling once before finetuning or in an on-policy form by periodically resampling during training. $N$ can be adjusted based on the available computational budget.
BS-T and BS-S are compatible with existing unlearning objectives such as NPO and WGA, and can also integrate regularization like GradDiff. As shown in §6, both bring clear gains: BS-T offers higher efficiency, while BS-S achieves more thorough forgetting. Pseudocodes are provided in Appx. C.
5 Theoretical Analysis: How BS Mitigates the Squeezing Effect?
This section establishes a unified theoretical perspective on how BS mitigates the squeezing effect. §5.1 revisits the AKG learning dynamics framework and illustrates how BS-T reshapes the residual term that drives token-level probability shifts. §5.2 extends this analysis to off-policy BS-S, which aggregates BS-T residuals over a broader set of belief-aligned continuations. Taken together, these results reveal how BS spreads forgetting pressure across both local belief neighborhoods and broader sequence-level alternatives. Detailed proofs and discussions are deferred to Appx. D.
5.1 BS-T: Residual Reshaping in the AKG Framework
We next formalize the learning dynamics underlying BS-T. Our analysis builds on the learning dynamics framework of LLM finetuning (Ren and Sutherland, 2025), which characterizes how an SGD update on an unlearning pair $\bm{\mathsf{\chi}}_{\rm u}$ influences the log-probability of any candidate response ${\mathbf{y}}_{\rm o}$. This framework decomposes the update into three components: a softmax Jacobian $\mathcal{A}$ capturing normalization effects, a kernel term $\mathcal{K}$ transporting influence across examples, and a residual term $\mathcal{G}$ reflecting the direct action of the loss. Lem. 5.1 restates this decomposition and highlights the residual as the driver of probability shifts, and Thm. 5.2 compares the residuals of GA and BS-T to show how BS-T spreads forgetting pressure over both the target token and its local belief neighborhood.
Lemma 5.1 (AKG Decomposition (Ren and Sutherland, 2025)).
@@ 5.1 BS-T: Residual Reshaping in the AKG Framework
where $\mathcal{A}_{t}(\bm{\mathsf{\chi}}_{\rm o})=\bm{I}-\mathbbm{1}\pi^{\top}_{{\bm{\theta}}^{t}}(\cdot|\bm{\mathsf{\chi}}_{\rm o})$ is the softmax Jacobian, $\mathcal{K}_{t}(\bm{\mathsf{\chi}}_{\rm o},\bm{\mathsf{\chi}}_{\rm u})=\nabla_{{\bm{\theta}}}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o})\nabla_{{\bm{\theta}}}^{\top}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm u})$ is the eNTK, and $\mathcal{G}_{t}(\bm{\mathsf{\chi}}_{\rm u})=\nabla_{{\mathbf{z}}}\mathcal{L}(\bm{\mathsf{\chi}}_{\rm u})$ captures the residual term induced solely by the unlearning loss. Here ${\mathbf{z}}=h_{{\bm{\theta}}}(\bm{\mathsf{\chi}})$ denotes the token–logit matrix and all quantities are evaluated at ${\bm{\theta}}^{t}$.
Lem. 5.1 indicates that the update is mainly governed by $\mathcal{G}$: it determines which tokens are pushed down or up before being modulated by $\mathcal{A}$ and transported via $\mathcal{K}$. Therefore, distinguishing the different forgetting behaviors of GA and BS-T reduces to analyzing the formulation of their residuals.
Theorem 5.2 (Residual Structure of GA vs. BS-T).
Under Lem. 5.1, denote ${\mathbf{q}}^{i}=\mathrm{sg}\big[\pi_{{\bm{\theta}}^{t}}(\cdot|\bm{\mathsf{\chi}}_{\rm u})\big|_{\mathcal{H}^{(i)}_{k}}\big$], the residual terms $\mathcal{G}$ for GA and BS-T at position $i$ are: (1) For GA, $\mathcal{G}^{i}_{\rm GA}=\pi_{{\bm{\theta}}^{t}}(\cdot|\bm{\mathsf{\chi}}_{\rm u})-{\mathbf{e}}_{y_{\rm u}^{i}}$; (2) For BS-T, $\mathcal{G}^{i}_{\rm BST}=\pi_{{\bm{\theta}}^{t}}(\cdot|\bm{\mathsf{\chi}}_{\rm u})-\big((1-\lambda){\mathbf{e}}_{y_{\rm u}^{i}}+\lambda\,{\mathbf{q}}^{i}\big)$. Hence for any component $v\neq y_{\rm u}^{i}$, we have
@@
Remark. Fig. 3 gives an intuitive illustration for Thm. 5.2. In GA, the gray curve $\pi_{\bm{\theta}_{\mathrm{o}}}$ shows the distribution before unlearning and the green curve $\pi_{{\bm{\theta}_{\mathrm{u}}}}$ after unlearning: the residual $\mathcal{G}_{\rm GA}$ pushes down the target $y_{\rm u}$ but reallocates mass to nearby high-likelihood regions, leading to semantically similar rephrasings. In BS-T, the shaded area marks the top-$k$ belief ${\mathbf{q}}^{i}$, and the residual $\mathcal{G}_{\rm BST}$ distributes repulsion across both the target and its close alternatives. The resulting blue curve suppresses the whole neighborhood rather than creating a new peak, reducing rephrasings and enabling more generalizable unlearning.
5.2 Off-Policy BS-S: Kernel-Weighted Residual Aggregation
We now extend the AKG framework to BS-S, using Thm. 5.3 to show that off-policy BS-S induces an update equal to a kernel-weighted sum of BS-T residuals computed over a fixed set of belief-aligned continuations. For each forget prompt ${\mathbf{x}}_{\rm u}$, we sample $N$ high-confidence continuations $\{\tilde{{\mathbf{y}}}_{\rm u}^{(j)}\}$ from a reference model before finetuning, and keep these responses fixed throughout finetuning. Let the original pair be $\bm{\mathsf{\chi}}_{\rm u}^{0}=[{\mathbf{x}}_{\rm u};{\mathbf{y}}_{\rm u}]$ and the auxiliary sequences be $\bm{\mathsf{\chi}}_{\rm u}^{j}=[{\mathbf{x}}_{\rm u};\tilde{{\mathbf{y}}}_{\rm u}^{(j)}]$, and define the weights as $\omega_{0}=1-\lambda_{\rm BSS}$ and $\omega_{j}=\lambda_{\rm BSS}/N$. With $\mathcal{L}_{\rm BST}$ as the underlying loss, off-policy BS-S corresponds to applying BS-T to the weighted set $\{\bm{\mathsf{\chi}}_{\rm u}^{m}\}_{m=0}^{N}$, yielding the following learning dynamics.
Theorem 5.3 (Learning Dynamics of Off-Policy BS-S).
Under Lem. 5.1 and the off-policy BS-S construction above, a single SGD step with learning rate $\eta$ on the off-policy BS-S loss
$\mathcal{L}_{\rm BSS}^{\textnormal{off}}({\bm{\theta}};\bm{\mathsf{\chi}}_{\rm u})=\sum\nolimits_{m=0}^{N}\omega_{m}\mathcal{L}_{\rm BST}({\bm{\theta}};\bm{\mathsf{\chi}}_{\rm u}^{m})$
updates the log-probability of any candidate response ${\mathbf{y}}_{\rm o}$ on $\bm{\mathsf{\chi}}_{\rm o}=[{\mathbf{x}}_{\rm u};{\mathbf{y}}_{\rm o}]$ by
$\Delta\log\pi_{t}({\mathbf{y}}_{\rm o}|\bm{\mathsf{\chi}}_{\rm o})=-\eta\mathcal{A}_{t}(\bm{\mathsf{\chi}}_{\rm o})\sum\nolimits_{m=0}^{N}\omega_{m}\mathcal{K}_{t}(\bm{\mathsf{\chi}}_{\rm o},\bm{\mathsf{\chi}}_{\rm u}^{m})\mathcal{G}_{{\rm BST},t}(\bm{\mathsf{\chi}}_{\rm u}^{m})+\mathcal{O}(\eta^{2}).$
Here $\mathcal{G}_{{\rm BST},t}(\bm{\mathsf{\chi}})$ is the BS-T residual, whose token-wise components coincide with $\mathcal{G}^{i}_{\rm BST}$ in Thm. 5.2.
Remark. Thm. 5.3 indicates that off-policy BS-S corresponds to applying BS-T to an expanded and fixed training set consisting of the original forget pair $\bm{\mathsf{\chi}}_{\rm u}^{0}$ together with a collection of auxiliary sequences $\{\bm{\mathsf{\chi}}_{\rm u}^{j}\}_{j=1}^{N}$ drawn from a frozen belief distribution. Under the AKG view, each sequence $\bm{\mathsf{\chi}}_{\rm u}^{m}$ contributes a BS-T residual $\mathcal{G}_{{\rm BST},t}(\bm{\mathsf{\chi}}_{\rm u}^{m})$ to the update on a test response $\bm{\mathsf{\chi}}_{\rm o}$, with its influence scaled by the kernel similarity $\mathcal{K}_{t}(\bm{\mathsf{\chi}}_{\rm o},\bm{\mathsf{\chi}}_{\rm u}^{m})$ and weight $\omega_{m}$. Compared with BS-T, which relies solely on the residual at $\bm{\mathsf{\chi}}_{\rm u}^{0}$, off-policy BS-S distributes forgetting pressure across a broader group of high-likelihood sequences, yielding smoother updates and more stable sequence-level unlearning.
Note that in on-policy BS-S, the auxiliary sequences are resampled from the model during finetuning and therefore depend on the evolving parameters $\bm{\theta}$}, which violates the teacher-forcing assumption required by the AKG framework. We discuss the implications of this limitation in Appx. D.4.
Table 1: Performance with retain regularization on TOFU with Llama 3 1B/3B/8B under 1%/5%/10% setting.
Llama 3.2 1B Llama 3.2 3B Llama 3.1 8B
@@ 5.2 Off-Policy BS-S: Kernel-Weighted Residual Aggregation
  • Note: Agg. is the harmonic mean of Mem. and Util.. Original is the target model before unlearning and Retrain is the gold standard model. ↑/↓ indicate larger/smaller values are preferable. The best and runner-up results are bolded and underlined.
@@ 6.1 Experimental Setup
Benchmarks, Baselines, and Models. We assess unlearning performance across three benchmarks: TOFU (Maini et al., 2024), MUSE (Shi et al., 2025), and WMDP (Li et al., 2024b). Our approach is compared with representative baselines from OpenUnlearning (Dorna et al., 2025) incorporating retain regularization, including GradDiff (Maini et al., 2024), NPO (Zhang et al., 2024), RMU (Li et al., 2024b), SimNPO (Fan et al., 2025), and WGA (Wang et al., 2025b). We adopt a variety of LLM families for unlearning, including Llama 2 (Touvron et al., 2023), Llama 3 (Grattafiori et al., 2024), and Zephyr (Tunstall et al., 2024). Specifically, for TOFU, we employ Llama 3.2 1B/3B-Instruct and Llama 3.1 8B-Instruct. For MUSE and WMDP, we use Llama 2 7B-Chat and Zephyr-7B-$\beta$, respectively.
@@
For further details and introductions of the experimental setup, please refer to Appx. E.
@@ 6.2 Experimental Results
Results on TOFU. Tab. 1 summarizes results on TOFU under 1%, 5%, and 10% forget settings. Across all model scales (Llama 3 1B/3B/8B), our bootstrapping methods achieve superior performance. In particular, BS-S consistently delivers the best aggregate and memorization scores (e.g., Agg. 0.58/0.60/0.60 at 5% and 0.57/0.50/0.49 at 1%), clearly surpassing NPO and RMU. BS-T also ranks second in most cases, balancing forgetting and retention—for instance, Agg. 0.55 at 5%–3B and 0.54 at 1%–1B—while retaining competitive utility with higher efficiency. These findings confirm that unlearning both targets and model beliefs enables BS-S to achieve the most thorough forgetting, with BS-T as a strong runner-up, validating the effectiveness of our framework on TOFU.
Table 2: Performance with retain regularization on WMDP with Zephyr-7B-$\beta$.
@@
Results on WMDP. Tab. 2 presents results on WMDP with Zephyr-7B-$\beta$. Recall that the forget score corresponds to QA accuracy, where values closer to 0.25 indicate more randomized responses and thus stronger unlearning. Both BS-T and BS-S achieve lower scores on Bio (0.26) and Cyber (0.28/0.27) compared with NPO (0.27/0.30) and RMU (0.29/0.27), while also attaining higher MMLU retention (0.52 and 0.54 vs. 0.44–0.48 for most baselines) except for RMU (0.55). Overall, BS-S delivers the best trade-off, reaching near-random forgetting accuracy while preserving more utility than most competing methods.
(a) BS-T Probability Dynamics
(b) BS-S Probability Dynamics
(c) LaaJ Evaluation on TOFU 10%
Figure 4: Log-probability dynamics under (a) BS-T and (b) BS-S, and (c) the LaaJ evaluation comparison after unlearning. Our methods monotonically suppress the target and high-likelihood probabilities and simultaneously achieve better LaaJ scores, indicating mitigation of both squeezing effects and spurious unlearning.
Analyzing Squeezing and Spurious Unlearning. To demonstrate the effectiveness of our methods for mitigating the squeezing effect and spurious unlearning, Fig. 4 jointly presents the probability dynamics of our BS and the LaaJ evaluation. In Fig. 4(a) and 4(b), BS-T and BS-S monotonically decrease the target log-probability and the high-likelihood neighbors, alleviating the squeezing effect. Fig. 4(c) further shows that BS-T and BS-S obtain higher Naturalness and Similarity than baselines, indicating that our framework mitigates spurious unlearning and preserves fluent. Here we use Gemini 2.5 Flash (Comanici et al., 2025) as the LLM judge with Llama 3.1 8B on TOFU 10%.
Additional Results in Appx. F. Owing to space limitations, further results are deferred to Appx. F. In addition to the content already been mentioned above, Appx. F.3 reports results on MUSE (-News and -Books); Appx. F.4 provides qualitative comparisons of unlearned responses across different unlearning methods; Appx. F.5 presents ablation studies covering hyperparameter analysis and the influence of different unlearning losses in BS-S; and Appx. F.6 reports training time comparisons.